OpenStudy (volleyballlover55):

Solve x^3 = 1 over 8. 1 over 2 ±1 over 2 1 over 4 ±1 over 4

OpenStudy (anonymous):

\[x^3=\frac{ 1 }{ 8 }=\frac{ 1 }{ 2^3 }=\left( \frac{ 1 }{ 2 } \right)^3,x=?\]

OpenStudy (volleyballlover55):

1/2? A?

OpenStudy (anonymous):

correct.

OpenStudy (volleyballlover55):

A square piece of cheese covers 8 square inches of a plate. What is the approximate length of one side of the piece of cheese? (Approximate to the nearest hundredth inch.) 2.81 inches 2.82 inches 2.83 inches 2.84 inches

OpenStudy (anonymous):